# Environment as Policy: Learning to Race in Unseen Tracks

### v1

@@ Abstract
Reinforcement learning (RL) has achieved outstanding success in complex robot control tasks, such as drone racing, where the RL agents have outperformed human champions in a known racing track. However, these agents fail in unseen track configurations, always requiring complete retraining when presented with new track layouts. This work aims to develop RL agents that generalize effectively to novel track configurations without retraining. The naïve solution of training directly on a diverse set of track layouts can overburden the agent, resulting in suboptimal policy learning as the increased complexity of the environment impairs the agent’s ability to learn to fly. To enhance the generalizability of the RL agent, we propose an adaptive environment-shaping framework that dynamically adjusts the training environment based on the agent’s performance. We achieve this by leveraging a secondary RL policy to design environments that strike a balance between being challenging and achievable, allowing the agent to adapt and improve progressively. Using our adaptive environment shaping, one single racing policy efficiently learns to race in diverse challenging tracks. Experimental results validated in both simulation and the real world show that our method enables drones to successfully fly complex and unseen race tracks, outperforming existing environment-shaping techniques. Website: rpg.ifi.uzh.ch/env_as_policy/index.html.
@@ I Introduction
Reinforcement learning (RL) involves agents learning through trial and error by interacting with a pre-defined environment and maximizing the rewards based on these interactions. It has proven highly effective in various robotic control applications, demonstrating remarkable task performance across scenarios like dexterous manipulation [1, 2, 3], quadrupedal locomotion [4, 5], and agile quadrotor flight [6, 7, 8]. Previous research has shown that RL can even outperform human champions in the task of drone racing [9, 10], where an RL agent flies a quadcopter drone through a complex track at high speed, requiring quick decision-making and precise control.
@@ II Related Works
Reinforcement Learning for Robotics Reinforcement Learning (RL) has been extensively applied in robotics to facilitate continuous control in complex and dynamic environments. Recent studies have achieved remarkable success in various robotic domains. These include quadrotor control [9, 10, 17, 8, 18], legged locomotion [19, 5], and manipulation [20, 21], showcasing significant advancements in the field.
@@ III-A Racing Policy Training
The autonomous racing task can be framed as an optimization problem, where the objective is to minimize the time it takes for an agile quadrotor to pass through a predefined sequence of gates [22], as illustrated in Fig. 2. In this task, we define the observations as $\bm{o}_{\mathrm{racing}}=\left[\bm{p},\tilde{\bm{R}},\bm{v},\bm{\omega},a_{\text{prev}},\delta\bm{p}_{1},\delta\bm{p}_{2}\right]$, where $\bm{p}\in\mathbb{R}^{3}$ denotes the drone’s position, $\tilde{\bm{R}}\in\mathbb{R}^{6}$ is a vector comprising the first two columns of $\bm{R}_{WB}$ [32], $\bm{v}\in\mathbb{R}^{3}$ and $\bm{\omega}\in\mathbb{R}^{3}$ denote the linear and angular velocity of the drone, $a_{\text{prev}}$ represents the previous action from the actor policy, and $\delta\bm{p}_{1},\delta\bm{p}_{2}\in\mathbb{R}^{12}$ represent the relative difference in position of the four next gate corners ($4\times 3$) in the world frame. Here, $\delta\bm{p}_{1}$ represents the difference of the four corners of the next gate to pass between the current quadrotor position, and $\delta\bm{p}_{2}$ represents the positional difference of the corners between the next gate to pass and the gate after the next gate to pass on the race track. The total reward at time $t$, denoted as $r_{t}$, consists of several components:

### v2

@@ Abstract
Reinforcement learning (RL) has achieved outstanding success in complex robot control tasks, such as drone racing, where the RL agents have outperformed human champions in a known racing track. However, these agents fail in unseen track configurations, always requiring complete retraining when presented with new track layouts. This work aims to develop RL agents that generalize effectively to novel track configurations without retraining. The naïve solution of training directly on a diverse set of track layouts can overburden the agent, resulting in suboptimal policy learning as the increased complexity of the environment impairs the agent’s ability to learn to fly. To enhance the generalizability of the RL agent, we propose an adaptive environment-shaping framework that dynamically adjusts the training environment based on the agent’s performance. We achieve this by leveraging a secondary RL policy to design environments that strike a balance between being challenging and achievable, allowing the agent to adapt and improve progressively. Using our adaptive environment shaping, one single racing policy efficiently learns to race in diverse challenging tracks. Experimental results validated in both simulation and the real world show that our method enables drones to successfully fly complex and unseen race tracks, outperforming existing environment-shaping techniques. Website: http://rpg.ifi.uzh.ch/env_as_policy.
@@ I Introduction
Reinforcement learning (RL) involves agents learning through trial and error by interacting with a pre-defined environment and maximizing the rewards based on these interactions. It has proven highly effective in various robotic control applications, demonstrating remarkable task performance across scenarios like dexterous manipulation [1, 2, 3, 4], quadrupedal locomotion [5, 6], and agile quadrotor flight [7, 8, 9, 10, 11]. Previous research has shown that RL can even outperform human champions in the task of drone racing [12, 13], where an RL agent flies a quadcopter drone through a complex track at high speed, requiring quick decision-making and precise control.
@@ II Related Works
Reinforcement Learning for Robotics Reinforcement Learning (RL) has been extensively applied in robotics to facilitate continuous control in complex and dynamic environments. Recent studies have achieved remarkable success in various robotic domains. These include quadrotor control [12, 13, 20, 9, 21, 22], legged locomotion [23, 6], and manipulation [24, 25], showcasing significant advancements in the field.
@@ III-A Racing Policy Training
The autonomous racing task can be framed as an optimization problem, where the objective is to minimize the time it takes for an agile quadrotor to pass through a predefined sequence of gates [26], as illustrated in Fig. 2. In this task, we define the observations as $\bm{o}_{\mathrm{racing}}=\left[\tilde{\bm{R}},\bm{v},\bm{\omega},a_{\text{prev}},\delta\bm{p}_{1},\delta\bm{p}_{2}\right]$, where $\tilde{\bm{R}}\in\mathbb{R}^{6}$ is a vector comprising the first two columns of $\bm{R}_{WB}$ [36], $\bm{v}\in\mathbb{R}^{3}$ and $\bm{\omega}\in\mathbb{R}^{3}$ denote the linear and angular velocity of the drone, $a_{\text{prev}}$ represents the previous action from the actor policy, and $\delta\bm{p}_{1},\delta\bm{p}_{2}\in\mathbb{R}^{12}$ represent the relative difference in position of the four next gate corners ($4\times 3$) in the world frame. Here, $\delta\bm{p}_{1}$ represents the difference of the four corners of the next gate to pass between the current quadrotor position, and $\delta\bm{p}_{2}$ represents the positional difference of the corners between the next gate to pass and the gate after the next gate to pass on the race track. The total reward at time $t$, denoted as $r_{t}$, consists of several components: